2.) Consider the following circuit. The white wire represents a switch. That can move to the right and connect the capacitorplease answer all parts
Answer #1

a. From 0 to 3 seconds ,The capacitor will get charged by the batteru through 100 ohm resistor. After 3 sec when swicth is moved to other position, then the charged capacitor will behave as a battery and will discharge through 500 ohm resistor.

b. Charging equation is

Vc= Vo (1 -e^-t/RC)

Vc=voltage on capacitor, vo=supply voltage, t=time, R =resistance, C= capacitace

Vc= 5(1 -e^-3/100*200*10^-6)  

Vc= 5 volts, So after three second the capacitor is fully charged and behaves as on open switch.

c. Now lets use discharging equation

Vc= Vo e^-t/RC

Vc=5 *e^-3/500*200*10^-6

Vc= 4.7*10^-13 volts, if you roound it to two three decimal points , its zero. i.e. its comletely dicharged.

d. For graph , First draw exponentially increasing graph,starting from zero volatage when time is zero and With max voltage is 5v and corresponding time is 3 seconds. and after three second exponentially decreasing graph with voltage becoming almost zero at t=6 seconds.